OpenStudy (anonymous):

What is the equation of the line which includes points (1, 5) and (-2, -7)? y=-4x-1 y=-4x+1/2 y=4x+1 y=-1/4x+1

OpenStudy (ghazi):

(y-5)=(-7-5)/(-2-1)* (x-1)

OpenStudy (ghazi):

this is called as two point slope form equation

OpenStudy (anonymous):

soo b?

OpenStudy (ghazi):

no

OpenStudy (ghazi):

y-5=4(x-1)

OpenStudy (anonymous):

c/D

OpenStudy (anonymous):

?

OpenStudy (ghazi):

y-5=4x-4 y=4x+1

OpenStudy (ghazi):

does it look like C

OpenStudy (anonymous):

c..

OpenStudy (anonymous):

thank youu!

OpenStudy (ghazi):

:)